Two trains leave from the same city, one going west, one going east, and one train is 15 mph faster than the other. After two hours they are 150 miles apart, how fast is each train going?
Answer by richwmiller(17219)

2*(x+x+15)=150
2*(2x+15)=150
4x+30=150
4x=120
x-30 mph
x+15=45 mph